# Is Vibe Coding Safe? Benchmarking Vulnerability of Agent-Generated Code in Real-World Tasks

### v1

@@ Abstract
Vibe coding is a new programming paradigm in which human engineers instruct large language model (LLM) agents to complete complex coding tasks with little supervision. Although it is increasingly adopted, are vibe coding outputs really safe to deploy in production? To answer this question, we propose SusVibes, a benchmark consisting of 200 feature-request software engineering tasks from real-world open-source projects, which, when given to human programmers, led to vulnerable implementations. We evaluate multiple widely used coding agents with frontier models on this benchmark. Disturbingly, all agents perform poorly in terms of software security. Although 61% of the solutions from SWE-Agent with Claude 4 Sonnet are functionally correct, only 10.5% are secure. Further experiments demonstrate that preliminary security strategies, such as augmenting the feature request with vulnerability hints, cannot mitigate these security issues. Our findings raise serious concerns about the widespread adoption of vibe-coding, particularly in security-sensitive applications.
Figure 1: SusVibes example task: An agent is started inside a docker environment and tasked with adding a feature to an existing code base. It take actions, interacts with the environment, collects feedback, and finally generates a patch to the repository. The generated solution patch is tested with unit tests targeting correctness and security. Without the line that calls check_unsafe_options, the patch cannot pass the security tests.
@@ 1 Introduction
Vibe coding is a new programming paradigm in which software engineers give natural language requests of software tasks and large language model (LLM) agents follow them to complete complicated programming tasks. It has been increasingly adopted, as indicated by the popularity of AI-based integrated development environments like Cursor and command-line interfaces like Claude Code. A recent survey shows that 75% of respondents are using vibe coding, among which 90% find it satisfactory [perry2025vibe]. Another survey suggests that beginner programmers with less than a year’s experience are much more likely to be vibe coding optimists [wired_ai_coding_2025]. Frontier AI companies, such as Anthropic, admittedly use “vibe coding in prod[uction]” [anthropic2024vibe]. While vibe coding may have increased engineer productivity, the security of agent-generated code remains questionable, especially when vibe coding users may not have the ability or intent to examine it carefully. Various sources report security incidents such as API keys being as plaintext and authentication vulnerabilities, some of which have already been exploited by malicious parties [archibald_kaplan_2025_vibe_coding].
@@
To address these limitations, we propose SusVibes, a benchmark to examine the security risks of AI agents for vibe coding. SusVibes consists of 200 realistic coding tasks on large repositories instead of single files and covers a wide range of 77 weaknesses from Common Weakness Enumeration (CWE) [CWE_mitre]. Its tasks are more complex, requiring editing on average 180 lines of code spanning multiple files. Table 1 compares SusVibes with existing secure code generation benchmarks, while Figure 1 shows an example task framed as requesting a feature (a unit of functionality that satisfies a requirement) [article] for an existing repository. An agent under evaluation is required to generate a patch to the repository that adds this feature. The patch is then tested with two sets of human-written unit tests, one for functional correctness and the other for security.
We propose an automatic pipeline that constructs SusVibes tasks from real-world GitHub repositories that contain fixed security issues. From the version history of a repository with a human-fixed vulnerability, we collect tests that were used to examine the vulnerability of a feature (e.g., a function) as security tests. We collect unit tests for the feature from the commit prior to the fix as functional correctness tests. We use the version of the repository two commits before the fix with the feature elided as the initial context of the task, and construct the task description (feature request) using LLM.
@@
Baxbench [vero2025baxbench] 392 (27) none N/A 13
CWEval [peng2025cweval] 119 file 10 31
SALLM [siddiq2024sallm] 100 file 12.9 45
SecCodePLT [yang2024seccodeplt] 1337 function 8.1 27
Asleep [pearce2025asleep] 89 file 19.6 18
ASE [lian2025ase] 120 repository 35.7 4
@@
We evaluate 3 open-source agent frameworks on top of 3 LLMs on SusVibes, resulting in 6 combinations in total. Disturbingly, even though the best-performing model, Claude 4 Sonnet, is able to solve 61.0% of the tasks and pass functional tests, over $80\%$ of its functionally correct solutions have vulnerabilities, exposing them to malicious exploitation. The specific problems solved securely are largely distinct across methods. Stratifying by vulnerability types (CWEs) shows that different frontier LLMs or frameworks favor different categories, leaving complementary strengths and blind spots.
Furthermore, we examine several preliminary attempts to mitigate security risks through prompting strategies, including adding generic security guidance (generic), using prompting to identify the CWE risk (self-selection), and providing the oracle CWE that this task targeted as a reference (oracle). Nevertheless, we find that although these strategies can improve the code security, the functional correctness is dropped significantly (about 6% points). This is because the agent focuses more on the security checks, making it pay less attention to the functionality it requires to implement. This leads to a drop in the number of overall correctly and securely solved tasks and calls for a more advanced vulnerability mitigation strategy in agent scenarios.
@@
  • We conduct a comprehensive set of experiments, which show that frontier LLMs and popular software engineering agents, despite their great ability to solve more than 50% of tasks and pass functional tests, perform very poorly in security, failing over 80% of security tests.
@@ 2 Related Work
Heralded by rapidly increasing performance on SWE-Bench [jimenezswe], LLM coding agents have become a big success in software engineering. Coding agents — LLM-based systems that take actions and interact with coding projects — can perform various tasks, including bug fixing, feature implementation, test generation [mundler2024swt], environment setup [eliseeva2025envbenchbenchmarkautomatedenvironment], or even generating a whole library from scratch [zhaocommit0].
Improvements for coding agents fall into two categories: agent design and model training. The former studies how to improve the agent scaffolding around the LLM: what actions are available to an agent [yang2024sweagent], what workflow an agent should follow [xia2025demystifying], how an agent can spend more inference-time compute in trade for better performance [antoniadesswe, zhangdiversity, gao2025trae]. The latter studies how to train a better LLM, supporting the agent. SWE-Gym [pan2024training] and SWESynInfer [ma2024lingma] train a single model for the agent with supervised-finetuning. SWE-Fixer [xie2025swe], CoPatcheR [tang2025co], SWE-Reasoner [ma2025thinking] train specialized models for different aspects of the agent, reducing the size of the model needed to achieve good performance. SEAlign [zhang2025sealign], SoRFT [ma2025sorft], and SWE-RL [wei2025swe] use reinforcement learning to train the model with either direct preference optimization or test results as rewards.
Despite a great amount of efforts into improving the capabilities of coding agents, few have focused on benchmarking and improving their security. SusVibes gives the community a platform to work on in this direction.
@@
Various benchmarks have emerged to assess both the security and the correctness of the LLM-generated code. Earlier ones focus on evaluating single-turn model generation in smaller scopes, such as a single file or a single function. SALLM [siddiq2024sallm] provides a framework to evaluate LLMs’ abilities to generate secure code with security-centric prompts. CWEval [peng2025cweval] introduces an outcome-driven evaluation framework that simultaneously assesses both functionality and security of LLM-generated code on the same problem set across multiple programming languages. SecCodePLT [yang2024seccodeplt] provides a unified platform for evaluating both insecure code generation and cyberattack helpfulness, combining expert-verified data with dynamic evaluation metrics in real-world attack scenarios. Asleep [pearce2025asleep] assesses the security of AI-generated code by investigating GitHub Copilot’s propensity to generate vulnerable code across three dimensions: diversity of weaknesses, prompts, and domains, finding approximately 40% of generated programs to be vulnerable.
More recent benchmarks have expanded their scope to repository level tasks with potential multi-file edits to be made. BaxBench [vero2025baxbench] focuses on backend application security by combining coding scenarios with popular backend frameworks across multiple programming languages, including functional and security test cases and expert-designed security exploits. ASE [lian2025ase] and SecureAgentBench [chen2025secureagentbenchbenchmarkingsecurecode] mines repository level vulnerability-fixing commits and repurpose them as tasks. Compared to these benchmarks, SusVibes focuses on evaluating coding agents, rather than models alone, covers significantly more CWEs and requires more lines to be edited. A detailed comparison between these secure code generation benchmarks is demonstrated in Table 1.
@@ 3 SusVibes: Developing a Security-Oriented Software Engineering Benchmark
Figure 2: Curation pipeline of mining open-source vulnerability commits, adaptively creating feature masks and task descriptions, and harnessing functionality and security tests. $\mathcal{C}_{0}$ is the vulnerability fixing commit, $\mathcal{C}_{-1}$ is the previous commit of $\mathcal{C}_{0}$, and $\mathcal{C}_{-1}^{\mathcal{M}}$ is the repository without feature implementation of $\mathcal{F}$. The security risk analysis in this example can be found in Section 4.3.
One common usage of vibe coding is specification to feature: the user provides some specification of a new feature and prompts an agent to implement the feature. When an inexperienced programmer overly relies on vibe coding to implement new features, it poses security risks, especially when the implementation shows plausible behavior. To mimic this use case, we present a method to automatically construct software engineering tasks aiming to expose the vulnerabilities of agent-implemented new features. These tasks are constructed from 108 existing open-source software projects across 10 security domains on GitHub. Each task corresponds to a historically observed security issue on a project. The agent’s solution could potentially touch many lines of code across multiple files. We also build environments to execute agent-generated solutions and evaluate their functional correctness and security. The resulting benchmark, SusVibes, contains 200 tasks covering 77 CWEs.
@@ 3.1 Benchmark Construction
The core principle of how a task in SusVibes is created is by selecting a commit $\mathcal{C}_{0}$ that fixes a known vulnerability in an existing feature $\mathcal{F}$ (verify_password()), reverting to the previous commit $\mathcal{C}_{-1}$ before the fix, and masking out $\mathcal{F}$ from its vulnerable implementation in $\mathcal{C}_{-1}$ to obtain $\mathcal{C}_{-1}^{\mathcal{M}}$. From this repository without $\mathcal{F}$, we create a task that requests the feature and harness tests for both functionality and security, as shown in Figure 2.
Harnessing Security Tests $\mathcal{T}_{secure}$ from Vulnerability Fixing Commits
We start by collecting over $20,000$ open-source, diverse vulnerability fixing commits over the last $10$ years from existing vulnerability fix datasets such as ReposVul [wang2024reposvulrepositorylevelhighqualityvulnerability] and MoreFixes [10.1145/3663533.3664036], yielding $\sim$3,000$\$ in Python. We focus on projects that use Python $\geq 3.7$ to avoid vulnerabilities tied to outdated versions and tooling dependencies. We further filter out the commits that do not modify the test suite, because those would not contain security tests that can detect the fixed vulnerabilities.
For a single vulnerability fixing commit $\mathcal{C}_{0}$, we separate the changes it made $\mathcal{P}$ into two parts — $\mathcal{P}^{\mathcal{F}}$ that modifies the implementation of $\mathcal{F}$ and $\mathcal{P}^{\mathcal{T}}$ that modifies the test suite, i.e. $\mathcal{P}=\mathcal{P}^{\mathcal{F}}+\mathcal{P}^{\mathcal{T}}$. In Figure 2, $\mathcal{P}^{\mathcal{F}}$ modifies secure_headers.py to fix a vulnerable feature implementation $\mathcal{F}$ (verify_password()), and $\mathcal{P}^{\mathcal{T}}$ modifies test_secure_headers.py which provides tests targeting the vulnerability (test_make_password_calls()). We use $\mathcal{P}^{\mathcal{F}}$ to locate the feature $\mathcal{F}$ that got fixed, and $\mathcal{P}^{\mathcal{T}}$ to collect added tests. The added tests from the vulnerability fixing commits are collected as the potential security tests $\mathcal{T}_{secure}$, and they can be added to the repository by applying $\mathcal{P}^{\mathcal{T}}$.
Harnessing $\mathcal{T}_{func}$ and Masking Out the Solution Code $\mathcal{F}$
Prompt I: Feature Masking $\mathcal{M}$ GOAL: Given a diff patch $\mathcal{P}^{\mathcal{F}}$, produce a deletion mask that removes a coherent implementation area enclosing this patch—i.e., delete all touched lines plus sufficient surrounding context. KEY DEFINITIONS:
@@
- Implementation area: enclosing logical units (e.g. function, class, block)
@@
- Use a tone like reporting Github issues
@@
Prompt III: Mask Verification GOAL: You are given a task description requesting a new feature, and a code patch $\mathcal{C}_{0}-\mathcal{C}_{-1}^{\mathcal{M}}$ purporting to implement it.
Your goal is to decide whether the patch contains any excessive implementation beyond what the task requires.
@@
Figure 3: SWE-agent is used to create the feature mask $\mathcal{M}$ (left), the task description (middle), and task verification (right), via operating on a software repository.
After harnessing $\mathcal{T}_{secure}$ from the vulnerability fixing commit $\mathcal{C}_{0}$, we checkout to the previous commit $\mathcal{C}_{-1}$, which contains the vulnerable implementation of $\mathcal{F}$, and the corresponding functionality tests $\mathcal{T}_{func}$. To synthesize a proper task from existing code, we utilize SWE-Agent [yang2024swe] to create a minimal mask that encloses the existing implementation of $\mathcal{F}$. SWE-Agent is started inside the code base at commit $\mathcal{C}_{-1}$, and given $\mathcal{P}^{\mathcal{F}}$, the unapplied modification to $\mathcal{F}$. It is instructed to mask out the feature that $\mathcal{P}^{\mathcal{F}}$ is modifying (left part of Figure 3). The mask is generated as a patch $\mathcal{M}$ and it only contains deletion of lines without addition. $\mathcal{M}$ is then applied to $\mathcal{C}_{-1}$ to obtain $\mathcal{C}_{-1}^{\mathcal{M}}$, the code base with solution code $\mathcal{F}$ masked out, as the initial context for a task in SusVibes.
Generating Task Description
@@
Adaptively Verifying the Mask
To ensure the feature request generated from $\mathcal{M}$ can cover the canonical feature implementation with security fixes, we verify the description line by line and adaptively modify the mask. As Figure 4 shows, this verification pipeline is detailed below.
To check if the generated feature request accounts for all lines in $\mathcal{C}_{0}-\mathcal{C}_{-1}^{\mathcal{M}}$, we use a third instance of SWE-Agent (right part of Figure 3) to link each line in $\mathcal{C}_{0}-\mathcal{C}_{-1}^{\mathcal{M}}$ to a requirement in the feature request. In case any implementation goes beyond what the description requires, we go back to the mask generation step to generate a larger mask . This loop is repeated adaptively until the generated request matches the canonical implementation.
@@
Building Execution Environment
We run SWE-Agent on each vulnerability fix commit $\mathcal{C}_{0}$ to build the execution environment for the repository and validate the test suite. In particular, the agent is provided with location of tests in $\mathcal{P}^{\mathcal{T}}$, as a hint on the core mandatory tests it should execute through in complex testing setups. We instruct it to consult, in order: the pre-existing container configurations, the CI/CD pipeline in .github/workflows, and other documentation for reproducing the testing workflow, and invoke docker commands to create a new Docker image with successful installation and testing steps. We employed LMs to synthesize test output parsers given multiple samples of test suite run results. The detailed process and the instructions can be found in Appendix A.3.
Execution-Based Test Case Validation
To rigorously validate tests for security and functionality based on execution results, we run different combinations of implementations and test suites, i.e. $\{\mathcal{C}_{0},\mathcal{C}_{-1},\mathcal{C}_{-1}^{\mathcal{M}}\}\times\{\mathcal{T}_{func},\mathcal{T}_{func}+\mathcal{T}_{secure}\}$. A valid task should satisfy the following requirements: (i) the masked vulnerable commit $\mathcal{C}_{-1}^{\mathcal{M}}$ must fail both functional and secure tests; (ii) the code base with vulnerable implementation $\mathcal{C}_{-1}$ needs to pass functional tests but fail secure tests; and (iii) the vulnerability fix commit $C_{0}$ needs to pass both test cases.

### v2

@@ Abstract
Vibe coding is a new programming paradigm in which human engineers instruct large language model (LLM) agents to complete complex coding tasks with little supervision. Although vibe coding is increasingly adopted, are its outputs really safe to deploy in production? To answer this question, we propose SusVibes, a benchmark consisting of 200 feature-request software engineering tasks from real-world open-source projects, which, when given to human programmers, led to vulnerable implementations. We evaluate multiple widely used coding agents with frontier models on this benchmark. Disturbingly, all agents perform poorly in terms of software security. Although 61% of the solutions from SWE-Agent with Claude 4 Sonnet are functionally correct, only 10.5% are secure. Further experiments demonstrate that preliminary security strategies, such as augmenting the feature request with vulnerability hints, cannot mitigate these security issues. Our findings raise serious concerns about the widespread adoption of vibe-coding, particularly in security-sensitive applications.
Figure 1: SusVibes: A feature-request task consists of a task description, a codebase, and an execution environment. The software engineering agent is asked to add the new feature to the given codebase. The agent can interact with the environment to get feedback and refine its solution patch. The solution patch will be tested with both human-written functionality and security unit tests. As the example shows, the solution patch cannot pass the security tests without check_unsafe_options.
@@ 1 Introduction
Vibe coding is a new programming paradigm in which software engineers give natural language requests of software tasks, and large language model (LLM) agents follow them to complete complicated programming tasks. It has been increasingly adopted, as indicated by the popularity of AI-based integrated development environments like Cursor and command-line interfaces like Claude Code. A recent survey shows that $75\%$ of respondents are using vibe coding, among which $90\%$ find it satisfactory [perry2025vibe]. Another survey suggests that beginner programmers with less than a year’s experience are much more likely to be vibe coding optimists [wired_ai_coding_2025]. Frontier AI companies, such as Anthropic, admittedly use “vibe coding in prod[uction]” [anthropic2024vibe]. While vibe coding may have increased engineer productivity, the security of agent-generated code remains questionable, especially when vibe coding users may not have the ability or intent to examine it carefully. Various sources report security incidents such as API keys being as plaintext and authentication vulnerabilities, some of which have already been exploited by malicious parties [archibald_kaplan_2025_vibe_coding].
@@
To address these limitations, we propose SusVibes, a benchmark to examine the security risks of AI agents for vibe coding. SusVibes consists of 200 realistic coding tasks on large repositories instead of single files and covers a wide range of 77 weaknesses from Common Weakness Enumeration (CWE) [CWE_mitre]. Its tasks are more complex, requiring editing on average 170 lines of code spanning multiple files. Table 1 compares SusVibes with existing secure code generation benchmarks, while Figure 1 shows an example task framed as requesting a feature (a unit of functionality that satisfies a requirement) [article] for an existing repository. An agent under evaluation is required to generate a patch to the repository that adds this feature. The patch is then tested with two sets of human-written unit tests, one for functional correctness and the other for security.
We propose an automatic pipeline that constructs SusVibes tasks from real-world GitHub repositories that contain fixed security issues. This pipeline includes three steps: (i) mining open-source repositories with human-fixed vulnerabilities; (ii) harnessing human-written functionality and security tests; (iii) adaptively generating the feature implementation mask, task description, and execution environment. We further propose a detailed verification phase to verify whether the generated feature request and execution environment are sufficient and necessary for feature implementation.
@@
Baxbench [vero2025baxbench] $392$ ($27$) none N/A $13$
CWEval [peng2025cweval] $119$ file $\times$ $10$ $31$
SALLM [siddiq2024sallm] $100$ file $\times$ $12.9$ $45$
SecCodePLT [yang2024seccodeplt] $1337$ function $\times$ $8.1$ $27$
Asleep [pearce2025asleep] $89$ file $\times$ $19.6$ $18$
ASE [lian2025ase] $120$ repository $\times$ $35.7$ 44
@@
We evaluate $3$ agent frameworks on top of $3$ LLMs on SusVibes, resulting in $9$ combinations. We find that even though the best-performing combination, Claude 4 Sonnet with SWE-Agent, is able to solve $61.0\%$ of the tasks and pass functional tests, over $80\%$ of its functionally correct solutions have vulnerabilities, exposing them to malicious exploitation. Results stratified by vulnerability types (CWEs) show that different frontier LLMs or frameworks favor different categories, leaving complementary strengths and blind spots.
Furthermore, we examine several preliminary attempts to mitigate security risks through prompting strategies, including adding generic security guidance (generic), using prompting to identify the CWE risk (self-selection), and providing the oracle CWE that this task targeted as a reference (oracle). However, while these strategies improve code security, they significantly reduce functional correctness by about $7$ percentage points. We hypothesize that this reduction arises because the agent prioritizes security checks, thereby paying less attention to implementing the required functionality. As a result, the preliminary mitigation decreases the number of tasks that are both functionally correct and secure, highlighting the need for more advanced vulnerability mitigation strategies in agent-based settings.
@@
  • We conduct a comprehensive set of experiments, which show that frontier LLMs and popular software engineering agents, despite their great ability to solve more than $50\%$ of tasks and pass functional tests, perform very poorly in security, failing over $80\%$ of security tests.
@@ 2 Related Work
Heralded by rapidly increasing performance on SWE-Bench [jimenezswe], LLM coding agents have become a big success in software engineering. Coding agents — LLM-based systems that take actions and interact with coding projects — can perform various tasks, including bug fixing, feature implementation, test generation [mundler2024swt], environment setup [eliseeva2025envbench], or even generating a whole library from scratch [zhaocommit0].
Improvements for coding agents fall into two categories: agent design and model training. The former studies how to improve the agent scaffolding around the LLM: what actions are available to an agent [yang2024swe], what workflow an agent should follow [xia2025demystifying], how an agent can spend more inference-time compute in trade for better performance [antoniadesswe, zhangdiversity, gao2025trae]. The latter studies how to train a better LLM, supporting the agent. SWE-Gym [pantraining] and SWESynInfer [ma2024lingma] train a single model for the agent with supervised-finetuning. SWE-Fixer [xie-etal-2025-swe], CoPatcheR [tang2025co], SWE-Reasoner [ma2025thinking] train specialized models for different aspects of the agent, reducing the size of the model needed to achieve good performance. SEAlign [zhang2025sealign], SoRFT [ma2025sorft], and SWE-RL [wei2025swerl] use reinforcement learning to train the model with either direct preference optimization or test results as rewards.
Despite a great amount of effort to improve the capabilities of coding agents, few have focused on benchmarking and improving their security. SusVibes gives the community a platform to work in this direction.
@@
Various benchmarks have emerged to assess both the security and the correctness of LLM-generated code. Earlier ones focus on evaluating single-turn model generation in smaller scopes, such as a single file or a single function. SALLM [siddiq2024sallm] provides a framework to evaluate LLMs’ abilities to generate secure code with security-centric prompts. CWEval [peng2025cweval] introduces an outcome-driven evaluation framework that simultaneously assesses both functionality and security of LLM-generated code on the same problem set across multiple programming languages. SecCodePLT [yang2024seccodeplt] provides a unified platform to evaluate both insecure code generation and cyberattack helpfulness, combining expert-verified data with dynamic evaluation metrics in real-world attack scenarios. Asleep [pearce2025asleep] assesses the security of AI-generated code by investigating the propensity of GitHub Copilot to generate vulnerable code across three dimensions: diversity of weaknesses, prompts, and domains, finding approximately 40% of generated programs to be vulnerable.
More recent benchmarks have expanded their scope to repository-level tasks with potential multi-file edits to be made. BaxBench [vero2025baxbench] focuses on backend application security by combining coding scenarios with popular backend frameworks across multiple programming languages, including functional and security test cases and expert-designed security exploits. ASE [lian2025ase] and SecureAgentBench [chen2025secureagentbenchbenchmarkingsecurecode] mines repository level vulnerability-fixing commits and repurpose them as tasks. Compared to these benchmarks, SusVibes focuses on evaluating coding agents, rather than models alone, covers significantly more CWEs and requires more lines to be edited. A detailed comparison between these secure code generation benchmarks is demonstrated in Table 1.
@@ 3 SusVibes: Developing a Security-Oriented Software Engineering Benchmark
Figure 2: SusVibes curation pipeline. It includes 33 steps: (i) mining open-source vulnerability fix commits $\mathcal{C}_{0}$ and reverting to its previous commit $\mathcal{C}_{0}$; (ii) harnessing human-written tests $\mathcal{T}_{func}$ and $\mathcal{T}_{secure}$; and (iii) adaptively creating feature masks and task descriptions. $\mathcal{C}_{-1}^{\mathcal{M}}$ is the repository without feature implementation of $\mathcal{F}$.
One common usage of vibe coding is specification to feature generation: a user gives natural language descriptions of a specific software functionality (i.e., a feature) based on an initial repository, and an agent autonomously generates code with various external tools such as a compiler. When an inexperienced programmer overly relies on vibe coding to implement new features, it poses security risks, especially when the implementation shows plausible behavior. Our approach will automatically construct software engineering tasks which expose the vulnerabilities of agent-implemented feature code. These tasks are sourced from $108$ existing open-source software projects across $10$ security domains on GitHub. Each task corresponds to a historically observed security issue on a project. The agent’s solution could potentially touch many lines of code in multiple files. We also build environments to execute agent-generated solutions and evaluate their functional correctness and security. The resulting SusVibes contains $200$ tasks covering $77$ CWEs.
@@ 3.1 Benchmark Construction
Our main idea to construct a task in SusVibes consists of three steps: selecting-reverting vulnerability fix commit, harnessing human-written software tests, and masking core code, as in Figure 2. First, we select a commit $\mathcal{C}_{0}$ from an existing software repository that fixes a known vulnerability in an existing feature $\mathcal{F}$ (e.g. verify_password ). We then revert to the preceding commit $\mathcal{C}_{-1}$ touching the requested functionality before the fix. Then, we compare the commits and identify the software tests for the feature’s functionality and security. Thirdly, we use an LLM to mask out the feature code $\mathcal{F}$ in $\mathcal{C}_{-1}$ to obtain $\mathcal{C}_{-1}^{\mathcal{M}}$. From this version without $\mathcal{F}$, we create a task requesting the feature.
Mining Open-source Vulnerability Fix Commits.
We start by collecting over $20,000$ open-source, diverse vulnerability fixing commits over the last $10$ years from existing vulnerability fix datasets, ReposVul [wang2024reposvul] and MoreFixes [10.1145/3663533.3664036], yielding $\sim$3,000$\$ in Python. We focus on software projects that use Python $\geq 3.7$ to avoid vulnerabilities tied to outdated versions and tooling dependencies. We further filter out the commits that do not modify tests, because those would not contain security tests that can detect the fixed vulnerabilities. More details about the mining process are in Appendix A.1.
Harnessing Security Tests $\mathcal{T}_{secure}$ and $\mathcal{T}_{func}$.
For a single vulnerability fixing commit $\mathcal{C}_{0}$, we separate the changes it made $\mathcal{P}$ into two parts — $\mathcal{P}_{\mathcal{F}}$ that modifies the implementation of $\mathcal{F}$ and $\mathcal{P}_{\mathcal{T}}$ that modifies the test suite, i.e. $\mathcal{P}=\mathcal{P}_{\mathcal{F}}+\mathcal{P}_{\mathcal{T}}$.In Figure 2, $\mathcal{P}_{\mathcal{F}}$ modifies secure_headers.py to fix a vulnerable feature implementation $\mathcal{F}$ (verify_password), and $\mathcal{P}_{\mathcal{T}}$ modifies test_secure_headers.py which provides tests targeting the vulnerability (test_make_password_calls). We use $\mathcal{P}_{\mathcal{F}}$ to locate the feature $\mathcal{F}$ that got fixed, and $\mathcal{P}_{\mathcal{T}}$ to collect added tests. The added tests from vulnerability fixing commits are collected as potential security tests $\mathcal{T}_{secure}$, they can be added to the repository by applying $\mathcal{P}_{\mathcal{T}}$. After harnessing $\mathcal{T}_{secure}$ from the vulnerability fixing commit $\mathcal{C}_{0}$, we checkout to the previous commit $\mathcal{C}_{-1}$, which contains the vulnerable implementation of $\mathcal{F}$, and the corresponding functionality tests $\mathcal{T}_{func}$.
Prompt I: Feature Masking $\mathcal{M}$ GOAL: Given a diff patch $\mathcal{P}_{\mathcal{F}}$, produce a deletion mask that removes a coherent implementation area enclosing this patch—i.e., delete all touched lines plus sufficient surrounding context. KEY DEFINITIONS:
@@
- Implementation area: enclosing logical units (e.g. func, class, etc.)
@@
- Use a GitHub issue tone
@@
Prompt III: Mask Verification GOAL: You’re given a task description requesting a new feature, and a code patch $\mathcal{C}_{0}-\mathcal{C}_{-1}^{\mathcal{M}}$ purporting to implement it.
You need to decide whether the patch contains any excessive implementation beyond what the task requires.
@@
Figure 3: Prompts for creating the feature mask $\mathcal{M}$ (left), generating the task description (middle), and verifying the mask (right); full versions are provided in Appendix A.2.
Generating Solution Code Mask $\mathcal{F}$ and Task Description.
To synthesize a proper task from existing code, we utilize SWE-agent [yang2024swe] to create a minimal mask that encloses the existing implementation of $\mathcal{F}$. SWE-agent is started inside the code base at commit $\mathcal{C}_{-1}$, and given $\mathcal{P}_{\mathcal{F}}$, the unapplied modification to $\mathcal{F}$. It is instructed to mask out the feature that $\mathcal{P}_{\mathcal{F}}$ is modifying (left part of Figure 3). The mask is generated as a patch $\mathcal{M}$ and it only contains deletion of lines without addition. $\mathcal{M}$ is then applied to $\mathcal{C}_{-1}$ to obtain $\mathcal{C}_{-1}^{\mathcal{M}}$, the code base with solution code $\mathcal{F}$ masked out, as the initial context for a task in SusVibes.
@@
3.2 Task Verification and Execution Environment
Adaptively Verifying Mask.
To ensure the generated feature request generated from $\mathcal{M}$ can cover the canonical feature implementation with security fixes, we verify the description line by line and adaptively modify the mask. To check if the generated feature request accounts for all lines in $\mathcal{C}_{0}-\mathcal{C}_{-1}^{\mathcal{M}}$, we use a third instance of SWE-agent (right part of Figure 3) to link each line in $\mathcal{C}_{0}-\mathcal{C}_{-1}^{\mathcal{M}}$ to a requirement in the feature request. In case any implementation goes beyond what the description requires, we go back to the mask generation step to generate a larger mask. This loop is repeated adaptively until the generated request matches the canonical implementation. Figure 4 illustrates the verification workflow.
Creating Execution Environment.
We run SWE-agent on each vulnerability fix commit $\mathcal{C}_{0}$ to build the execution environment for the repository and validate the test suite. In particular, the agent is provided with the location of tests in $\mathcal{P}_{\mathcal{T}}$, as a hint on the core mandatory tests it should execute through in complex testing setups. We instruct it to consult, in order: the pre-existing container configurations, the CI/CD pipeline in .github/workflows, and other documentation for reproducing the testing workflow, and invoke docker commands to create a new Docker image with successful installation and testing steps. The detailed process and the instructions can be found in Appendix A.3.
Validating Test Cases via Execution.
To validate tests for security and functionality based on execution results, we run different combinations of implementations and test suites, i.e. $\{\mathcal{C}_{0},\mathcal{C}_{-1},\mathcal{C}_{-1}^{\mathcal{M}}\}\times\{\mathcal{T}_{func},\mathcal{T}_{func}+\mathcal{T}_{secure}\}$. A valid task should satisfy the following requirements: (i) the masked vulnerable commit $\mathcal{C}_{-1}^{\mathcal{M}}$ must fail both functional and secure tests; (ii) the code base with vulnerable implementation $\mathcal{C}_{-1}$ needs to pass functional tests but fail secure tests; and (iii) the vulnerability fix commit $C_{0}$ needs to pass both test cases.